OpenStudy (anonymous):

Multiply. (x-2)(x+5) Simplify your answer.

OpenStudy (anonymous):

you need to foil it out

OpenStudy (anonymous):

look up the graph if you type it in it will give you the answer

OpenStudy (anonymous):

there will be a graph that shows up and it will give you your answer

OpenStudy (anonymous):

when we foil we multiply the first, the outside, the inside, and the last

OpenStudy (anonymous):

then we get (x-2)(x-5)=x^2-5x-2x+7

OpenStudy (anonymous):

x squared-3x-10

OpenStudy (anonymous):

then we simplify to get x^2-7x+7

OpenStudy (anonymous):

well x^2-3x-10

OpenStudy (anonymous):

this is the correct answer

OpenStudy (anonymous):

Draw a line (in your head or on paper) from x to x, -5 to 2, -5 to x, and x and 2. Now we need to multiply these figures. So basically multiply everything that's in the brackets. So now we have, X x X= x squared -5x2= -10 -5 x X= -5x X x 2=2x Add the x's together, 2x-5x= -3x We now have -10, x squared and -3x so... (x - 5)(x + 2)= x squared-3x-10

OpenStudy (anonymous):

haha my bad ya when we foil we x^2-3x+10 my bad

OpenStudy (anonymous):

I fiqured out the answer its x^2+3x-10

OpenStudy (anonymous):

there you go

OpenStudy (anonymous):

good job